### v2

# KamLAND-Zen Collaboration

Abstract
We present a search for neutrinoless double-beta ($0\nu\beta\beta$) decay of 136Xe using the full KamLAND-Zen 800 dataset with 745 kg of enriched xenon, corresponding to an exposure of $2.1$ ton yr of 136Xe. This updated search benefits from a more than twofold increase in exposure, recovery of photo-sensor gain, and reduced background from muon-induced spallation of xenon. Combining with the search in the previous KamLAND-Zen phase, we obtain a lower limit for the $0\nu\beta\beta$ decay half-life of $T_{1/2}^{0\nu}>3.8\times 10^{26}$ yr at 90% C.L., a factor of 1.7 improvement over the previous limit. The corresponding upper limits on the effective Majorana neutrino mass are in the range 28–122 meV using phenomenological nuclear matrix element calculations.
The search for neutrinoless double-beta ($0\nu\beta\beta$) decay is an active field of research in nuclear and particle physics Agostini2023 . Observation of this decay would prove the Majorana nature of neutrinos and demonstrate the non-conservation of lepton number, a key feature of Leptogenesis, which can explain the imbalance of matter over antimatter in the Universe Fukugita1986 . In the model of light Majorana neutrino exchange, the $0\nu\beta\beta$ decay rate is proportional to the square of the effective Majorana neutrino mass $\left<m_{\beta\beta}\right>\equiv\left|\Sigma_{i}U_{ei}^{2}m_{\nu_{i}}\right|$. The most stringent limit to date has been provided by KamLAND-Zen, reaching the inverted mass ordering (IO) region below 50 meV for the first time Abe2023a . An improved search covering the unexplored IO region would stringently test theoretical models predicting $\left<m_{\beta\beta}\right$> in this range Harigaya2012 ; Asaka2020 ; Asai2020 , with a significantly increased chance to observe $0\nu\beta\beta$ decay Agostini2017 .
[…]
KamLAND-Zen 800completed data-taking in January 2024 to start dismantling the current KamLAND detector in preparation for a detector upgrade, KamLAND2-Zen. This Letter reports on the $0\nu\beta\beta$ decay analysis with the full KamLAND-Zen 800 dataset collected between February 5, 2019, and January 12, 2024, and its combination with the earlier KamLAND-Zen 400 data.
[…]
During the earlier period in KamLAND-Zen 800, we observed an increasing trend in the number of low photo-detection efficiency PMTs caused by age-related gain decrease. The resulting degradation in energy resolution could increase the background from the high energy tail of $2\nu\beta\beta$ events. In March 2020, we started a continuous campaign modifying the input circuitry of the KamLAND front-end electronics (FEE) to counter the PMT gain decrease. The 1 p.e. gain and timing offsets due to the circuit modification are corrected for each PMT based on photon counting with high-statistics background events. Owing to these restoration efforts, we have succeeded in preventing the degradation of the energy resolution by a factor of 1.2, which would have led to a twofold increase of the $2\nu\beta\beta$ background.
Figure 1: Vertex distribution of candidate singles data (SD) events (black points) in the $0\nu\beta\beta$ energy window $2.35<E<2.70\,{\rm MeV}$. The thick black and red lines indicate the shape of the IB and the 1.57-m-radius spherical volume, respectively. The dot-dashed line indicates the nylon belt suspending the IB. The thin dashed lines illustrate the shape of the equal-volume spherical half-shells, which compose the 2.5-m-radius spherical fiducial volume. The high-count region at the IB bottom within the yellow solid line indicates the hot spot and is removed through a 0.7 m radius cut.
We apply the following cuts to select $\beta\beta$ decay events: (i) The reconstructed vertex must be within 2.5 m of the detector center and 0.7 m away from the bottom hot spot on the IB, which is outlined in Fig. 1. (ii) Muons and events within 2 ms after muons are rejected. (iii) Bi-Po decay pairs are eliminated by a delayed coincidence tag, which requires the time and distance between the prompt Bi and delayed Po decay candidates to be less than 1.9 ms and 1.7 m, respectively, and by rejecting events with a double pulse identification within a single event acquisition window. Those cuts remove $(99.95\pm 0.03)$% of 214Bi-214Po events, and $(97.9\pm 0.5)$% of 212Bi-212Po. (iv) Reactor $\overline{\nu}_{e}$’s identified by a delayed coincidence of positrons and neutron-capture $\gamma$’s are rejected. (v) Poorly reconstructed events are rejected. These events are tagged using a vertex-time-charge discriminator which measures how well the observed PMT time-charge distributions match those expected based on the reconstructed vertex. The overall selection inefficiency for $0\nu\beta\beta$ events is less than 0.1%.
Background sources external to the Xe-LS are dominated by radioactive impurities (RI) on the IB film. The contamination levels of 238U and 232Th are $(4.3\pm 0.2)\times 10^{-12}$ g/g and $(2.4\pm 0.1)\times 10^{-11}$ g/g, respectively, which are roughly a factor of 10 smaller than those measured on the previous IB Gando2016 . The reference calculations for 238U and 232Th mentioned here assume secular equilibrium for comparison with the previously reported values. In the earlier period of the dataset, we found an increase in the background rate at the IB bottom, possibly due to the settling of dust particles containing radioactive impurities. To avoid this possible background, we tag and remove this high-background period from the dataset using machine learning algorithms, as discussed in Ref. Abe2023a ; Li2023 .
Solar neutrinos are an intrinsic background source for KamLAND-Zen. Based on neutrino flux calculations using the standard solar model Serenelli2011 , the elastic scattering (ES) of 8B solar neutrinos on electrons in the Xe-LS is estimated to be $(4.9\pm 0.2)\times 10^{-3}$ (ton day)-1, including the effect of three-flavor neutrino oscillations. The charged-current (CC) interactions on 136Xe produce $e^{-}$, 136Cs, and $\gamma$’s from the excited states, and the subsequent decays of 136Cs ($\tau=19.0$ days, $Q=2.548$ MeV) create a background peak around 2.0 MeV in visible energy, mostly overlapping with the resolution tail of $2\nu\beta\beta$ decays. The interaction rate is expected to be $(0.8\pm 0.1)\times 10^{-3}$ (ton day)-1 based on the cross section calculated in Ref. Ejiri2014 ; Frekers2013 .
Cosmic-ray muons produce neutrons and radioactive isotopes through nuclear spallation of carbon and xenon, which decay by emitting $\beta$’s or $\gamma$’s. To remove these events, we apply cuts based on space and time correlation with muon and neutron capture $\gamma$ events. In the $0\nu\beta\beta$ window, decays of 10C ($\tau=27.8$ s, $Q=3.65$ MeV) and 6He ($\tau=1.16\,{\rm s}$, $Q=3.51\,{\rm MeV}$) dominate the muon spallation backgrounds. To reduce these short-lived backgrounds, we remove events within 150 ms after muons, and events reconstructed within 1.6 m of neutron vertices for 180 s. In addition, we reject remaining background events by the “shower” likelihood method based on reconstructed muon shower profiles Li2014 ; Li2015a ; Li2015b ; Zhang2016 ; Abe2023b . The overall rejection efficiencies for 10C and 6He are >>99.799.7% and $(97.3\pm 1.5)\%$, respectively. To reduce the 137Xe ($\tau=5.5\,{\rm min}$, $Q=4.17\,{\rm MeV}$) background, we remove events reconstructed within 1.6 m of vertices identified as neutron captures on 136Xe, which produce high energy $\gamma$’s ($Q=4.03$ MeV), for 27 min. This cut removes $(74\pm 7)\%$ of 137Xe. The deadtime introduced by these cuts is $(14.6\pm 0.1)\%$.
The products of muon spallation with lifetimes greater than $\mathcal{O}$(100 s), denoted as “long-lived products”, are attributed to the decay of heavy isotopes produced by xenon spallation. Xenon spallation can be tagged by detecting multiple neutrons. To characterize the long-lived products, we define a likelihood ratio, $R_{L}=L_{\rm spa}/(L_{\rm spa}+L_{\rm acc})$. Here $L_{\rm spa}$ and $L_{\rm acc}$ are the probability density functions (PDFs) for long-lived muon-spallation pairs and accidental pairs, respectively. These PDFs are constructed as a function of neutron multiplicity, distance to neutron vertices, and time interval from preceding muons. The cut value on $R_{L}$ is optimized using an estimate with MC simulation tools. FLUKA BOHLEN2014211 ; Ferrari:898301 is used to calculate the spallation isotope yields, and Geant4 to calculate the radioactive decays and energy spectra of the isotopes, considering their sequential decay chain. The predicted backgrounds are primarily from 132I, 130I, 124I, 122I, 118Sb, 110In, 88Y, and from other isotopes in smaller amounts (Table IX in Ref. Abe2023b ). The total background rate in the $0\nu\beta\beta$ window is $30.0\pm 2.2$ (ton yr)-1 in total, in units of 136Xe exposure. Events that are not classified as coming from long-lived backgrounds are referred to as “singles data” (SD), and the others are referred to as “long-lived data” (LD).
Figure 2: Energy spectra of selected $\beta\beta$ candidates within a 1.57-m-radius spherical volume drawn together with best-fit backgrounds, the $2\nu\beta\beta$ decay spectrum, and the 90% C.L. upper limit for $0\nu\beta\beta$ decay of (a) singles data (SD), and (b) long-lived data (LD). The LD exposure is about 10% of the SD exposure.
We improved the neutron identification for post-muon events to enhance the tagging efficiency of the long-lived backgrounds. We identify these neutrons with a second, deadtime-free electronics system, called MoGURA Abe2023b . Previously, the neutron vertex was reconstructed using only PMTs with normal gain. In this analysis, we added low gain PMTs to the photon counting, achieving an increase in the mean number of PMT hits from $181$ to $201$ for 2.225 MeV $\gamma$’s from neutron capture, after subtracting afterpulse noise contributions. The resulting neutron tagging efficiency is estimated to be $(74.5\pm 0.4)\%$ based on the neutron capture time distribution, approximately 2% better than in the previous analysis. Our MC study shows that $(44.0\pm 8.7)\%$ of long-lived spallation backgrounds are classified as LD, whereas only 8.98.9% of uncorrelated events are mis-classified.
The total livetime for SD and LD in KamLAND-Zen 800 is $1133$ days and $111$ days, respectively. This is the effective measurement time considering the space and time cuts described in the background section. The total exposure of SD, which is sensitive to the $0\nu\beta\beta$ signal, is $2.1$ ton yr of 136Xe, more than double the amount of the previous search Abe2023a . The $0\nu\beta\beta$ decay rate is estimated from a simultaneous likelihood fit (Statistics section in Ref. PDG2024 ) to the binned energy spectra of SD and LD between 0.5 and 4.8 MeV, where systematic uncertainties are profiled in the likelihood. The binned energy spectra are made by dividing the 2.5-m-radius fiducial volume into 40 equal-volume bins, as illustrated in Fig. 1. The time variation of event rates is also considered in the fit. The contributions from major backgrounds in the Xe-LS, such as 85Kr, 40K, 210Bi, the 228Th-208Pb subchain of the 232Th series, and long-lived spallation products, are free parameters and are left unconstrained in the spectral fit. The contributions from the  [^2] Rn-210Pb subchain of the 238U series and short-lived spallation products are constrained by their independent measurements Abe2010 . The parameters of the detector energy response model are constrained by 214Bi data. For the long-lived spallation products an energy spectral distortion parameter based on linear rescaling is applied and constrained by its estimated uncertainty (Fig. 10 in Ref. [26]).
Figure 3: Allowed region for the 136Xe $0\nu\beta\beta$ rate and the long-lived spallation background rate in the energy region $2.35<E<2.70\,{\rm MeV}$ ($0\nu\beta\beta$ window). The red dot and contour lines correspond to the best-fit and the $1\sigma$, 90%, 95%, 99% C.L. intervals for two degrees of freedom, respectively. The horizontal line indicates the MC-based prediction.
Figure 2 shows the energy spectra of selected candidate SD and LD events within a 1.57-m-radius spherical volume together with the best-fit curves. The exposure of 136Xe for SD in this reduced volume is 1.13 ton yr. The best-fit background contributions are summarized in Table 1. The $0\nu\beta\beta$ signal best-fit is 0 events, indicating no event excess over the background expectation. Based on a profile likelihood method, we obtain a 90% confidence level (C.L.) upper limit on the number of $0\nu\beta\beta$ decays of <<$9.7$ events (<<$7.8$ events in the range $2.35<E<2.70\,{\rm MeV}$), which corresponds to a limit of <<$8.8$ (ton yr)-1 in units of 136Xe exposure, or $T_{1/2}^{0\nu\beta\beta}>3.5\times 10^{26}$ yr (90% C.L.). An analysis based on the Feldman-Cousins procedure PhysRevD.57.3873_FCmethod gives a slightly stronger limit of $4.3\times 10^{26}$ yr (90% C.L.), indicating that the physical boundary at zero has only a limited impact on the $0\nu\beta\beta$ rate estimate, even with low statistics. An MC simulation of an ensemble of experiments assuming the best-fit background spectrum without a $0\nu\beta\beta$ signal indicates a median sensitivity of $2.2\times 10^{26}$ yr. The probability of obtaining a limit stronger than that reported here is 20%. In addition to the frequentist analyses above, we also performed a statistical analysis within the Bayesian framework, assuming a flat prior for $1/T_{1/2}^{0\nu\beta\beta}$. The Bayesian limit and median sensitivity are $3.4\times 10^{26}$ yr and $2.4\times 10^{26}$ yr (90% C.L.), respectively.
Figure 4: Effective Majorana neutrino mass $\left<m_{\beta\beta}\right$> as a function of the lightest neutrino mass $m_{\rm lightest}$. The dark shaded regions are predictions based on best-fit values of neutrino oscillation parameters for (a) the normal ordering (NO) and (b) the inverted ordering (IO), and the light shaded regions indicate the $3\sigma$ ranges calculated from oscillation parameter uncertainties DellOro2014 ; NuFIT2020 . The horizontal lines indicate 90% C.L. upper limits on $\left<m_{\beta\beta}\right$> with 136Xe from KamLAND-Zen (this work), considering an improved phase space factor calculation Kotila2012 ; Stoica2013 and phenomenological NME calculations: shell model Menendez2018 ; PhysRevC.93.024308horoi ; PhysRevC.101.044315.coraggio2020 ; PhysRevC.105.034312.coraggio2022 (dot-dashed lines), quasiparticle randomphase approximation (QRPA) PhysRevC.87.064302.mustonen ; PhysRevC.91.024613.Hyv ; PhysRevC.98.064325.simko2018 ; PhysRevC.97.045503fang ; PhysRevC.102.044303terasaki (dotted lines), energy-density functional (EDF) theory PhysRevLett.105.252503.rodriguez ; PhysRevLett.111.142501lopez ; PhysRevC.95.024305song (solid lines), interacting boson model (IBM) PhysRevC.91.034304barea ; PhysRevD.102.095016deppisch (dashed lines). (c) The corresponding limits for 136Xe, 76Ge Agostini2020 , and 130Te Adams2022 . Three theoretical predictions in the IO region are also shown: (A)Harigaya2012 , (B)Asaka2020 , (C)Asai2020 .
Table 1: Summary of the estimated and best-fit background contributions in the energy region $2.35<E<2.70\,{\rm MeV}$ within the 1.57-m-radius spherical volume. In total, $60$ events were observed.
[h] Background        Estimated       Best-fit 136Xe $2\nu\beta\beta$ $27.6\pm 0.2$ $\dagger$ $27.68$ Residual radioactivity in Xe-LS 238U series $\pm 0.01 0.080.08$ 232Th series - $1.35$ External (Radioactivity in IB) 238U series - $6.61$ 232Th series - $0.04$ Neutrino interactions 8B solar $\nu$ $e^{-}$ ES $3.57\pm 0.09$ $3.58$ Spallation products Long-lived $17.82\pm 1.30$ $\dagger$ $20.27$ 10C $\pm 0.05 0.000.00$ 6He $0.54\pm 0.27$ $0.56$ 137Xe $0.72\pm 0.60$ $0.71$
  • $\dagger$
    Estimations based on the previous half-life measurements (136Xe $2\nu\beta\beta$) and the spallation MC study (long-lived). Those event rate constraints are not applied in the spectral fit.
Figure 3 shows the allowed region of the 136Xe $0\nu\beta\beta$ rate and the long-lived spallation background rate from a combined fit of the KamLAND-Zen 400 and 800 datasets, giving a limit of
$T_{1/2}^{0\nu\beta\beta}>3.8\times 10^{26}\,\textrm{yr~~~(90\% C.L.)}.$ (1)
The best-fit long-lived spallation background rate is $30.2^{+5.0}_{-4.0}$ (ton yr)-1indicating good consistency between the MC-based prediction and the LD analysis. This combined analysis has a sensitivity of $2.4\times 10^{26}$ yr, and the probability of obtaining a stronger limit is 2020%. The lower limit on $T_{1/2}^{0\nu\beta\beta}$ can be converted to an upper limit on $\langle m_{\beta\beta}\rangle$ assuming that the $0\nu\beta\beta$ decay mechanism is dominated by exchange of light Majorana neutrinos. Assuming the axial coupling constant $g_{A}\simeq 1.27$, we obtain an upper limit of
$\langle m_{\beta\beta}\rangle<(28-122)\,\textrm{meV~~~(90\% C.L.)},$ (2)
using an improved phase space factor calculation of $14.54\times 10^{-15}\,(\textrm{yr})^{-1}$ Kotila2012 ; Stoica2013 , and phenomenological nuclear matrix element (NME) calculations, summarized in Table 2. In this letter, we follow the current standard to use those phenomenological models, although recent theoretical work has produced new NME calculations belley2023ab ; Jokiniemi2023 incorporating both quenching physics Faessler_2008 ; PhysRevC.87.014315 and short-range contact PhysRevLett.120.202001 contributions, which are missing in the calculations in Table 2. Considering the calculations from Ref. belley2023ab ; Jokiniemi2023 gives ranges of upper limits on $\langle m_{\beta\beta}\rangle$ of 7112571-125 meV (ab initio belley2023ab ), 256025-60 meV (pnQRPA Jokiniemi2023 ), and 4413844-138 meV (shell model Jokiniemi2023 ). Figure 4 illustrates the allowed range of $\langle m_{\beta\beta}\rangle$ as a function of the lightest neutrino mass for the NME values in Table 2. The shaded regions include the uncertainties in mixing elements $U_{ei}$ and the neutrino mass splitting for each hierarchy. Also drawn are experimental limits from $0\nu\beta\beta$ decay searches in other nuclei Agostini2020 ; Adams2022 . The limit using 136Xe by KamLAND-Zen has provided the most stringent tests of the effective Majorana mass in the IO region and several theoretical models Harigaya2012 ; Asaka2020 ; Asai2020 . It also provides the strongest constraint on $m_{\rm lightest}$ considering extreme cases of the combination of CP phases and the uncertainties from neutrino oscillation parameters DellOro2014 ; NuFIT2020 . We obtain a 90% C.L. upper limit of $m_{\rm lightest}<(84-353)\,{\rm meV}$.
In conclusion, we presented a $0\nu\beta\beta$ decay search with the complete KamLAND-Zen dataset including improvements in the rejection of muon-induced long-lived spallation backgrounds. This search places the most stringent constraints to date on $\left<m_{\beta\beta}\right$>, and reaches the inverted mass ordering region below 50 meV with half of the phenomenological nuclear matrix element calculations. The sensitivity is limited by background from xenon spallation products and $2\nu\beta\beta$ decay. In the near future, we plan to upgrade the detector (KamLAND2-Zen) with more 136Xe mass, enhanced energy resolution to minimize the leakage of the $2\nu\beta\beta$ tail into the $0\nu\beta\beta$ analysis window, and enhanced rejection of spallation products through improved neutron tagging and particle identification methods.
Table 2: Summary of commonly used phenomenological nuclear matrix elements ($M^{0\nu}$) for 136Xe $0\nu\beta\beta$ decay and the corresponding 90% C.L. upper limits on the effective Majorana neutrino mass $\left<m_{\beta\beta}\right$> from KamLAND-Zen data. The limits shown in Eq. (2) use the minimum and maximum.
     Ref.    $M^{0\nu}$    $\langle m_{\beta\beta}\rangle$ (meV)
Shell model Menendez2018 2.28, 2.45 59.4, 55.3
PhysRevC.93.024308horoi 1.63, 1.76 83.1, 77.0
PhysRevC.101.044315.coraggio2020 ; PhysRevC.105.034312.coraggio2022 2.39 56.7
QRPA PhysRevC.87.064302.mustonen 1.55 87.4
PhysRevC.91.024613.Hyv 2.91 46.6
PhysRevC.98.064325.simko2018 2.71 50.0
PhysRevC.97.045503fang 1.11, 1.18 122, 115
PhysRevC.102.044303terasaki 3.38 40.1
EDF theory PhysRevLett.105.252503.rodriguez 4.20 32.3
PhysRevLett.111.142501lopez 4.77 28.4
PhysRevC.95.024305song 4.24 32.0
IBM PhysRevC.91.034304barea 3.25 41.7
PhysRevD.102.095016deppisch 3.40 39.9
Acknowledgements.
The KamLAND-Zen experiment is supported by JSPS KAKENHI Grants No. 21000001, No. 26104002, and No. 19H05803; the U.S. National Science Foundation awards no. 2110720 and no. 2012964; the Heising-Simons Foundation; the Dutch Research Council (NWO); and under the U.S. Department of Energy (DOE) Grant No. DE-AC02-05CH11231, as well as other DOE and NSF grants to individual institutions. The Kamioka Mining and Smelting Company has provided service for activities in the mine. We acknowledge the support of NII for SINET4.